OpenStudy (patrickcharran97):

In set builder notation, how do you write the solutions of the inequality? x-2 is less than or equal to -1?

OpenStudy (anonymous):

\[x-2 \leq -1\] \[x \leq 1\] \[ \left\{ - \infty,1\right\}\]

OpenStudy (anonymous):

\[(-\infty,1]\]